### v3

# Conversational SimulMT: Efficient Simultaneous Translation with Large Language Models

Abstract
Simultaneous machine translation (SimulMT) presents a challenging trade-off between translation quality and latency. Recent studies have shown that LLMs can achieve good performance in SimulMT tasks. However, this often comes at the expense of high inference cost and latency. In this paper, we propose a conversational SimulMT framework to enhance the inference efficiency of LLM-based SimulMT through multi-turn-dialogue-based decoding where source and target chunks interleave in translation history, enabling the reuse of Key-Value cache. To adapt LLMs to the proposed conversational decoding, we create supervised fine-tuning training data by segmenting parallel sentences using an alignment tool and a novel augmentation technique to enhance generalization. Our experiments with Llama2-7b-chat on two SimulMT benchmarks demonstrate that the proposed method empowers the superiority of LLM in translation quality, meanwhile achieves comparable computational latency with specialized SimulMT models. [^1] Code, weights, and data will be released with publication.
Conversational SimulMT: Efficient Simultaneous Translation
with Large Language Models
Minghan Wang1, Thuy-Trang Vu1, Yuxia Wang2, Ehsan Shareghi1, Gholamreza Haffari1 1Department of Data Science & AI, Monash University  2MBZUAI {minghan.wang,trang.vu1,ehsan.shareghi,gholamreza.haffari}@monash.edu yuxia.wang@mbzuai.ac.ae
1 Introduction
Simultaneous machine translation (SimulMT) systems provide real-time translation of text input stream Gu et al. (2017). This task plays an important role in real-world applications, such as facilitating communication in online conferences and generating live subtitles with strict latency requirements. In online speech translation, where re-translation is impossible, SimulMT should be robust to errors propagated from the automatic speech recognition system, and produce high-quality translation with low latency.
(a) Offline Prompt
(b) Conversational Prompt
Figure 1: SimulMT prompts in Llama2 format. The system message is blue. The source and corresponding target translations chunks have matching color codes.
Although large language models (LLMs) have shown the potentials in machine translation (Hendy et al., 2023; Zhu et al., 2023), their applications to SimulMT is non-trivial, as they are not inherently designed for simultaneous decoding. Recent works have attempted to adapt LLMs for SimulMT with prefix fine-tuning, incremental decoding (Wang et al., 2023b) and learning to wait for more source tokens before translation (Koshkin et al., 2024). These works show LLMs, with careful prompt-engineering, could approach the performance of specialized SimulMT models. However, high computational cost, slow inference, and high latency render these approaches impractical for real-world applications (Yuan et al., 2024). This is primarily due to the use of offline prompting, where arriving source tokens are inserted at the end of the source sequence, disrupting the continuity of the translation history (Figure 1(a)). This prevents reusing cached target history states and requires re-computation of source and target representations.
To mitigate this issue, we propose conversational prompts that resemble the multi-turn dialogue nature of LLMs. Specifically, user inputs are treated as the source tokens to be read, while the LLM’s responses are considered the predicted target tokens to be written. In our conversational SimulMT, newly arrived source form the current instruction, while previous source tokens and their translation are treated as conversation history (Figure 1(b)). This conversational prompt format enables the reuse of Key-Value cache Pope et al. (2023), as all content is appended incrementally without modifying the translation history, hence improving inference efficiency. However, conversational SimulMT poses a new challenge for LLMs to accurately identify source sequences, particularly when these sequences are interrupted with target chunks within the conversation history.
Interleaving incomplete source and target segments in the dialogue history is unnatural (see Figure 1(b)). The code-switching phenomenon is exhibited to some extent in LLMs for some languages (Yong et al., 2023); however, it is the continuation rather than the translation of the previous content, making it challenging to leverage existing code-switched datasets for training. To adapt the LLM to the conversational decoding format, we take the approach of supervised fine-tuning (SFT) in the conversational SimulMT. The challenge is the lack of the conversational SimulMT data for SFT. We generate the training data by segmenting parallel sentence pairs into smaller chunks based on a transformation of the word alignments. The segmented chunks are further augmented to handle different latency requirements in SimulMT.
Experiments on two SimulMT benchmarks demonstrate the effectiveness of our proposed conversational SimulMT in balancing the trade-offs between accuracy, speed and flexibility to different latency requirements. Compared to offline prompting, our method not only maintains strong performance, but also benefits from reduced latency. Notably, our method attains similar decoding speed to the offline LLM-based MT.
Our contributions are summarized as follows,
  • We introduce the conversation prompt to reduce the inference cost of LLM-based SimulMT by leveraging its multi-turn dialogue capability and enabling efficient reuse of Key-Value cached computations.
  • We present an automated training data curation pipeline which involves a simple heuristic and a novel augmentation strategy to efficiently extract source and target chunks from their word alignments, enabling the LLM to decode with prefix-growing code-switched inputs to support processing of incoming source words while generating target words.
  • Experiments demonstrate that the proposed conversational SimulMT obtains up to 2$\times$ acceleration compared to the offline-prompting baseline while maintaining comparable translation quality, emphasizing its value in practical applications.
2 Background
Simultaneous Machine Translation (SimulMT)
Unlike offline machine translation, where models generate target translation $\mathbf{y}=(y_{1},...,y_{J})$ given a complete source sentence $\mathbf{x}=(x_{1},...,x_{I})$, SimulMT incrementally translates in an online manner with partial source context $\mathbf{x}_{\leq t}=(x_{1},...,x_{t})$ where $I$ and $J$ are source and target length, and $t\leq I$. A core component of SimulMT is a read-write policy to decide at every step whether to wait for one or multiple new source tokens (READ) or to generate target tokens given the current partial source sentence (WRITE). This policy is crucial in balancing the trade-offs between translation quality and latency.
Incremental Decoding
Studies have explored equipping the capability of simultaneous decoding to an offline neural machine translation model by repetitively performing offline decoding on the incrementally updated source and target histories, without the necessity of READ -WRITE policy learning and model customization (Liu et al., 2020; Nguyen et al., 2021; Polák et al., 2022; Guo et al., 2023). Specifically, it involves a chunk-wise READ policy that reads a fixed number of $n$ tokens in each round and a WRITE policy that selects the longest common prefix (LCP) among the beam candidates as a stable hypothesis to be committed Polák et al. (2022). LCP often results in high latency when no candidates share the same initial token, hindering the WRITE action. Relaxed Agreement Longest Common Prefix (RALCP) (Wang et al., 2023b) employs a voting mechanism to accept a common prefix if the agreement among candidates is above an agreement threshold $\gamma$. Once a hypothesis is committed, it is no longer modified.
SimulMT with LLMs
Given that incremental decoding essentially repeats performing offline decoding, leveraging offline-style translation prompts for LLMs is a straightforward approach. Meanwhile, offline-style prompts align with LLMs’ innate ability to follow instructions (translating content from one language to another), avoiding the additional fine-tuning Xu et al. (2023). During each incremental decoding round, a new source chunk is READ in and appended to the source history $\mathbf{x}$. The translation hypothesis is WRITTEN out and appended to target history $\mathbf{y}$, forming the latest incremental states. Given the new source and target history $\mathbf{x}$ and $\mathbf{y}$, LLMs are prompted to generate the latest translation with the offline prompt as Figure 1(a).
3 Conversational SimulMT
While incremental decoding with offline prompt enables LLMs to perform simultaneous decoding, it faces high computational latency due to the insertion of newly arrived source tokens in the middle of prompt, disrupting the reuse of cached target history states. In this section, we propose conversational prompts to improve the decoding efficiency and balance quality-latency trade-off.
3.1 Decoding with Conversational Prompt
The efficiency improvement in LLMs hinges on maintaining the Key-Value (KV-) cache reuse, i.e. the decoding process must consistently add new tokens at the end of the sequence without altering the middle elements. When LLMs are performing multi-turn dialogues, the prompt for each turn is composed of a user input and agent response, separated by [INST] and [/INST] tags: “<s>[INST] $X$ [/INST] $Y$</s>", and conversation histories are simply concatenated as the context Touvron et al. (2023). Drawing parallels to multi-turn dialogues in LLMs, SimulMT can also be viewed similarly, where user inputs are equivalent of READ action, and agent responses are equivalent of WRITE action. At round $t$, LLM reads a source context chunk $X_{t}$ and writes its translation $Y_{t}$: “<s>[INST] $X_{t}$ [/INST] $Y_{t}$</s>". The conversational prompt of SimulMT can be constructed by placing the source chunk as user input, and the model produces the translation as a response (see Figure 1(b)). The conversational prompt enables re-using KV-cache as all contents are appended incrementally. Our approach is also able to leverage the hypothesis selection strategy of RALCP (Wang et al., 2023b) to prune the unstable suffixes in each response. Algorithm 1 in Appendix B presents the detailed decoding process.
Setting N-Shot SacreBLEU COMET
Offline MT 0-Shot 30.99 84.95
Convers. SimulMT 0-Shot 7.14 58.76
Convers. SimulMT 5-Shot 13.51 69.03
Convers. SimulMT 0-Shot Failure Case
Chunk 1 Input: Die Flugdaten zeigten, dass das
Chunk 1 Response:
The flight data showed that the plane
was flying at an altitude of 35,000 feet.
Chunk 2 Input: Flugzeug auch bei einem zweiten
Chunk 2 Response: The plane was also flying during the second flight.
Reference Flight data showed the plane had to pull out of second
Table 1: Performance of zero-shot and few-shot conversational SimulMT (Convers.) with Llama2-7b-chat on the WMT15 De->En test set. Offline decoding result with the same model is presented (Offline MT) as the baseline. A failure case is also shown, in which the LLM tends to complete the sentence in the translation with hallucination (red) leading to a more serious failure in the second chunk.
We conduct a pilot experiment on the zero-shot and few-shot capability of LLMs in decoding with conversation prompts. Specifically, we evaluate the pretrained Llama2-7b-chat (Touvron et al., 2023) with zero- and five-shot settings on the WMT15 De->En test set with chunk size $n$ set as 5. As illustrated by Table 1, SimulMT with conversational prompts obtain inferior performance even when instructed with 5-shot prompt. The failure case example indicates that the LLM cannot accurately translate the content within the partial source sentence. Instead, it tends to complete the sentence in the translation with hallucinated content. This is because the LLM is trained on full sentences and, therefore, is not able to handle interleaved partial sentence translation in the multi-turn dialogue format. To this end, we propose to perform supervised fine-tuning (SFT) with the LLM on SimulMT data in a conversational format. Our approach to tackle the challenge of lacking the conversational SimulMT data for SFT is introduced in the following section.
Figure 2: The illustration of the data curating process. The first graph is obtained from fast_align, it is then modified into a monotonic dependency graph by adding additional edges. The Meta Trajectory can be derived by segmenting the monotonic dependency graph with minimal dependency (segment with the dotted line in graph 3). Finally, Policy Generalization is applied to augment the segmented graph with merge (red dotted lines will be removed) and shift (blue dotted lines are shifted) operations. Chunks in the trajectories derived from the third and fourth graphs are highlighted with different colors.
3.2 SFT on Conversational SimulMT Data
As conversational SimulMT data is not naturally available, we propose to synthesize READ / WRITE chunks by segmenting sentence pairs from parallel corpora. Inspired by Arthur et al. (2021) which generates the oracle policy from word alignments, we further extend the approach by carefully addressing the impact of word reordering and improving the generalizability of the oracle policy. Specifically, we first build monotonic dependency graph from the alignment of a sentence pair. We then segment the graph and convert these segments into READ / WRITE pairs, followed by augmentation to improve its generalization across various latency demands (Figure 2). The process is explained below.
Alignment Graph Generation
Given a sentence pair, we employ fastalign Dyer et al. (2013) to obtain word alignment between source and target tokens (Step 1 in Figure 2). The obtained alignment is a set $\mathcal{A}$ of pairs $(i,j)$ denoting the source token $x_{i}$ is aligned with its corresponding target token $y_{j}$. We define the sufficient source token set to generate a given target token $y_{j}$ as $\textbf{a}_{j}=\{i|(i,j)\in\mathcal{A},\forall i\in[0,I]\}$ of all source tokens needed to produce the target token $y_{j}$.
A source and target sentences have a monotonic translation relationship if the previous target tokens only aligned with the previous source tokens, i.e. $\forall j>k\quad\max(\textbf{a}_{j})\geq\max(\textbf{a}_{k})$ (Koehn et al., 2005; Ling et al., 2011). This condition ensures that the relative order of words is preserved between the source and target sentences. In that case, the optimal minimum-latency policy that retains sufficient source information is to produce the monotonic translation that follows the word order of the source sequence, i.e. WRITE the target token $y_{j}$ after reading the final source token $x_{\max(\textbf{a}_{j})}$ needed for its translation, and READ the next source tokens.
Monotonic Dependency Graph
The monotonic dependency between source and target sentences enables the effective implementation of an optimal READ /WRITE policy. However, in practice, translations often exhibit non-monotonic relationships, which require re-ordering to produce grammatically correct output in the target language, especially when the source and target languages have significantly different syntactic structures. To address this, we propose a heuristic method for constructing a monotonic dependency graph $\overrightarrow{\mathcal{A}}$ from the alignment set $\mathcal{A}$, such that the monotonic condition is met (Step 2 in Figure 2).
More specifically, for each target token $y_{j}$ which does not satisfy the monotonic condition $\max(\textbf{a}_{j})<\max(\textbf{a}_{j-1})$, we add a new edge from the last sufficient source token $x_{\max(\textbf{a}_{j-1})}$ to $y_{j}$ to eliminate the need of re-ordering. For example, in Figure 2, $y_{2}$ violates the monotonic condition as the last source token in its sufficient set $\textbf{a}_{2}=\{1\}$ comes after the last sufficient source token of the previous target $\textbf{a}_{1}=\{1,2\}$. Therefore, a new edge from $x_{2}$ to $y_{2}$ is added to the graph.
Meta Trajectory
We then segment the monotonic dependency graph and convert these segments into READ / WRITE pairs, representing the meta trajectory of the oracle policy with minimum latency (Step 3 in Figure 2). This segmentation involves examining each target token and identifying its exclusive corresponding source tokens with minimal dependency. Each subgraph $\overrightarrow{\mathcal{A}_{j}}$ corresponds with a READ / WRITE pair ($R_{j}$, $W_{j}$) where $W_{j}=\{y_{j}\}$ can be a single target token and $R_{j}=\{x_{i}|i\in\textbf{a}_{j}\setminus\textbf{a}_{j-1}\}$ is a set of new source tokens since the last target token. When multiple consecutive target tokens depend on the same source token, we combine their WRITE actions and assign the shared source token to $R_{j}=\{x_{i}\}$, forming $W_{j}=\{y_{j},...,y_{j+n}\}$ with these target tokens. This procedure generates a meta READ / WRITE trajectory for the oracle policy, denoted as $RW^{\star}=[(R_{1},W_{1}),...,(R_{C},W_{C})],C\leq I$, representing $C$ chunks.
Trajectory Augmentation
Since the meta trajectories are tailored for minimal latency, they may not generalize well to different lengths of the input chunk, corresponding to different levels of latency. To improve the LLM’s adaptability across various latency demands, we augment the meta-trajectory $RW^{\star}$ with a series of merge and shift operations (Step 4 in Figure 2). We first traverse $RW^{\star}$ and randomly merge $\delta$ consecutive READ and WRITE actions, forming new pairs $([R_{c},...,R_{c+\delta}],[W_{c},...,W_{c+\delta}])$, where []delimited-[][$\cdot$][ ⋅ ] is the string concatenation operation. Here, $\delta$ is a variable re-sampled from a uniform distribution $\mathcal{U}(\delta_{\min},\delta_{\max})$ where $\delta_{\min}$ and $\delta_{\max}$ are predefined hyperparameters.
Additionally, with a probability of $\beta$, we randomly shift a portion of tokens from a WRITE action $W_{c}$ to the next one $W_{c+1}$ in the merged trajectory. More specifically, we split $W_{c}$ at a proportion $\rho$ and transfer the latter part to the next pair, resulting in $(R_{c},W_{c}^{<\rho}),(R_{c+1},[W_{c}^{>\rho},W_{c+1}])$ where $\rho$ is sampled from $\mathcal{U}(\rho_{\min},0.9)$ where $\rho_{\min}$ is a hyperparameter. The proposed augmentation strategy not only enhances the LLM’s context conditioning ability but also suits the incremental decoding, where the end of the prediction is often truncated by the stable hypothesis selection algorithm.
The resulting trajectory is a sequence of READ / WRITE chunks where each READ chunk contains source text of differing lengths. We then format them in the conversational format to make the trajectory into a text sequence. During SFT, we apply cross-entropy loss only on target tokens within WRITE chunks without the shift operation.
4 Experiments
Trajectory Dimension De-En En-Vi
Meta-Trajectory #Chunk $10.69\pm 5.51$ $12.98\pm 8.07$
#SRC word/Chunk $1.74\pm 0.78$ $1.38\pm 0.38$
#TGT word/Chunk $1.79\pm 0.78$ $1.73\pm 0.49$
Aug-Trajectory #Chunk $2.74\pm 1.21$ $3.12\pm 1.57$
#SRC word/Chunk $7.01\pm 3.92$ $5.83\pm 2.76$
#TGT word/Chunk $7.18\pm 3.90$ $7.35\pm 3.49$
Table 2: This table presents the statistics of the conversational SimulMT data curated from the training set of both benchmarks, including mean and standard deviation of the number of chunks, number of source/target tokens per chunk in the meta- and augmented-trajectory respectively.
4.1 Datasets
WMT15 De->En
(4.5M training pairs) We employ newstest2013 (3000 pairs) for validation and utilize newstest2015 (2169 pairs) as our test dataset222www.statmt.org/wmt15/. For baseline models, we follow instructions in Zhang et al. (2022) to use BPE Sennrich et al. (2016) for tokenization with 32,000 merge operations, the vocabulary is shared between German and English.
IWSLT15 En->Vi
(133K training pairs) TED tst2012 (1553 pairs) and tst2013 (1268 pairs) are used as validation and test set333nlp.stanford.edu/projects/nmt/. We follow Ma et al. (2020); Zhang et al. (2022) to preprocess the data for baseline models, resulting in vocabulary sizes of 17K and 7.7K for English and Vietnamese.
Note that the above tokenization is only applicable for non-LLM baseline models. For LLM-based SimulMT, we use their provided vocabularies and tokenizers.
Conversational SimulMT Dataset
We employ fastalign Dyer et al. (2013) to obtain word alignment between source and target tokens. We randomly merge $\delta_{\min}=2$ to $\delta_{\max}=10$ consecutive segments of the meta trajectories. For shift operation, we set the value of $\beta$ and $\rho_{\min}$ to 0.5, i.e. with the probability of 0.5, we shift at least 50% portions of tokens in a target segment to the next one. Table 2 presents detailed statistical information about the dataset.
4.2 Evaluation Metrics.
We use SacreBLEU [^4] BLEU+nrefs:1+case:mixed+eff:no+tok:13a
+smooth:exp+version:2.3.1
 (Post, 2018), COMET [^5] https://huggingface.co/Unbabel/wmt22-cometkiwi-da (Rei et al., 2020) and word-level average lagging (AL) (Ma et al., 2019) to evaluate the quality and latency. To measure the computational-aware latency, we use the word wall time (WWT) metric (Wang et al., 2023b), which is the average wall time to predict a word on the same hardware.
4.3 Model Training
For LLM-based methods, we follow Wang et al. (2023b) to utilize Llama2-7b-chat Touvron et al. (2023) as the backbone for both language pairs. We conduct SFT with QLoRA Hu et al. (2022); Dettmers et al. (2023) for one epoch to reduce the computation. The QLoRA adapters are configured with $r=64$ and $\alpha=16$. We set the learning rate to 2e-4, batch size to 48, and employ 4-bit quantization. For all experiments, a single A100 GPU is used. For offline prompt baselines, we tune the model on the offline bitext corpus. The LLMs using conversational prompts are tuned on our curated conversational SimulMT data (§3.2).
4.4 Settings
We compare our proposed conversational SimulMT method against the following baselines
Offline NMT
Following (Zhang and Feng, 2022), we train vanilla Transformers Vaswani et al. (2017) of size small (48M) for En->Vi and big (300M) for De->En, and use beam size of 5 during inference.
SimulMT
We report the performance of SimulMT with (i) fixed policy Wait-$k$ Ma et al. (2019) and (ii) adaptive policy, including Wait-Info Zhang et al. (2022) and ITST Zhang and Feng (2022).
  • Wait-$k$ (Ma et al., 2019) propose to achieve simultaneous decoding with an encoder-decoder Transformer by first reading $k$ source tokens and then alternately reading and writing one token per time. During decoding, we set $k$ from 1 to 8 for De->En and 4 to 8 for En->Vi.
  • Wait-Info (Zhang et al., 2022) quantitatively measures the information of each token, ensuring that the information in the target remains below the received source information with the threshold $\mathcal{K}$. We set the value of $\mathcal{K}$ from 1 to 8 for both datasets during decoding.
  • ITST (Zhang and Feng, 2022) calculates the weight of information transferred from each source token to the current target token. Based on the total information received, it then determines whether to proceed with the translation of the target token. We evaluate with the threshold ranging from 0.1 to 0.7 for both datasets in the experiment.
LLM-based SimulMT
We compare our conversational prompt against LLM-based SimulMT with offline prompt Wang et al. (2023b). For both methods, we set $n$=[3,5,7,9,11,13] in the experiment. Both approaches utilize RALCP as the stable hypothesis selection method if using beam search decoding (beam=5). Additionally, we also report GREEDY selection method (decoding with greedy search without pruning its prediction with prefix selection strategy), which serves as a low-latency baseline. LLM-based offline NMT is also reported as the performance upper bound.
Notations
For space reason, we abbreviate experimental settings in the following fields: (i) decoding mode: offline (O) or simultaneous (S), (ii) model name: dedicated SimulMT models or LLMs (LM), (iii) prompt type: offline (OP) or conversational prompt (CP) used only for LLMs, (iv) hypothesis selection method (only for LLM-based SimulMT): RALCP or GREEDY. For example, our proposed method for simultaneous decoding is written as S:LM:CP:RALCP.
Figure 3: Translation quality and latency results on two benchmarks. For abbreviations, please refer to the Notations description (§4). Results are presented in three groups with different colors: (i) specialized MT and SimulMT methods (yellow), (ii) LLM-based offline-prompting SimulMT (blue), and (iii) LLM-based conversational SimulMT (red).
4.5 Results
Our preliminary result in Table 1 have shown that out-of-the-box LLMs struggle in zero-shot and few-shot conversational SimulMT. In our experiments, we focus on evaluating the impact of fine-tuning LLMs on curated data, with the goal of improving their performance in simultaneous translation using conversational prompts. Specifically, we aim to determine if fine-tuning can enable LLMs to handle conversational SimulMT more effectively, and to explore any trade-offs between latency reduction and translation quality.
Translation Quality
Figure 3 shows the BLEU and COMET scores, as well as the average latency on two benchmarks. Overall, LLM-base SimulMT (red and blue) outperforms specialized methods (yellow), with the gain up to 3 BLEU scores and 6 COMET scores. This suggests that leveraging prior knowledge in LLMs is advantageous for translation. Furthermore, given increased latency allowance, LLM-based SimulMT can approach the performance of offline NMT. Among LLM-based methods, given the same latency, conversational SimulMT (red) achieves comparable results to its offline-prompting counterparts (blue) in BLEU and even performs better if measured by COMET. This shows our approach does not negatively impact the translation capabilities of LLMs.
Translation Latency
Compared to offline-promting method (blue), conversational SimulMT (red) can effectively reduce translation latency (1.17 and 1.50 AL in average for De->En and En->Vi) in all latency-range. It even reaches the similar latency level of specialized SimulMT (yellow) in En->Vi translation. The use of RALCP in conversational SimulMT (S:LM:CP:RALCP) can further boost quality over greedy strategy (S:LM:CP:GREEDY), though at the cost of increased latency due to its more conservative behavior. Compared to other baseline approaches, conversational SimulMT obtains a better trade-off between latency and quality, as it fits a wider latency span with more consistent quality.
More importantly, conversational SimulMT (red) can maintain better quality at the low-latency level (AL<4) compared to specialized models (yellow), making it more superior in practical application where both high quality and low latency are required. In contrast, given the same decoding configuration, decoding with offline-prompt (blue) is unable to fit into the low-latency span, making its quality advantageous less competitive compared to specialized approaches (yellow).
5 Analysis
Figure 4: This figure presents the relationship between computational-aware latency in Word Wall Time and the quality in COMET score. For settings of simultaneous decoding, they are represented by circles with the size representing the variance in both metrics under different latency control parameters such as $n$. For offline settings, diamond marks are used. Colors and notations are consistent with the Figure 3, where our proposed approach is highlighted with bold text.
5.1 Decoding Speed
While AL is effective in algorithmically quantifying the delay between the translation and the source input, it does not account for the computational cost associated with the inference process, i.e. assuming unit decoding speed across all models. In real-world applications, the actual inference wall time is a critical factor that directly impacts user experience. A model may exhibit low latency in terms of AL but incur a high inference cost, resulting in substantial time latency. To address this, we also report decoding speed using the word wall time (WWT) metric, which measures the real inference time per word (§4.4).
Detailed result in WMT15 De->En is shown in Figure 4. Overall, specialized MT and SimulMT The combination of offline-prompting and RALCP (S:LM:OP:RALCP) is the slowest configuration, significantly lagging behind its offline setting (O:LM:OP), despite of obtaining a good translation quality. Conversely, when employing conversational prompting (S:LM:CP:RALCP), it can achieve comparable decoding speed with the offline model (O:LM:CP) while maintaining the translation quality level. Moreover, employing greedy hypothesis selection with the conversational prompt (S:LM:CP:GREEDY) achieves a comparable speed to specialized SimulMT (yellow) but with superior quality.
Figure 5: The performance of the model tested with different levels of trajectory augmentation strategy.
5.2 Effectiveness of Trajectory Augmentation
We conduct an ablation study to assess the impact of trajectory augmentation methods by supervised fine-tuning LLMs on data derived from (i) meta trajectories only, (ii) incorporating merge operations, and (iii) both merge and shift operations (§3.2). The fine-tuning hyper-parameters are kept constant, with only the training data varies. Figure 5 reports the performance of conversational SimulMT with RALCP selection strategy under different training data configurations. Interestingly, compared to the model trained solely on meta trajectories, trajectory augmentation does not only enhance translation performance but also reduces latency, with the merge operation playing a pivotal role. We speculate that without augmentation, the model may have difficulty predicting accurately under varying latency conditions. This leads to RACLP accepting less stable hypotheses, hence worsening overall latency.
5.3 Ability of Context Following
Figure 6: The model performance for translating the last source chunk w/wo context under different chunk sizes.
Performing SimulMT with conversational prompts requires the LLM’s proficiency in utilizing context information accurately. To further examine the context-following capability of the model tuned with our curated data, we assess its performance in translating the final chunk of the source text, both with and without preceding context. Specifically, we extract the inference history for each instance, dividing it into a source-target dialogue history as context and the final source chunk as the latest input. Subsequently, the fine-tuned LLM is tasked with generating the last translation chunk from the last source chunk, either with or without the translation history. The evaluation is done by computing the BLEU score on the concatenation of the last chunk to its original history. A consistent 2-point decrease in performance can be found when decoding without context for all chunk sizes. This underscores the LLM’s ability to effectively harness context information for producing more precise translations, as illustrated in Figure 6.
6 Conclusion
This paper focuses on the feasibility of utilizing LLM for SimulMT. We found that leveraging the incremental-decoding framework with offline prompting leads to high computational latency, hindering the reuse of the Key-Value cache. To address this, we propose the conversational prompting which allows LLMs to conduct SimulMT in a multi-turn dialogue manner. The approach significantly speeds up the inference and also preserves the quality superiority, enabling practical LLM-based SimulMT systems.
Limitations
We summarize the limitations of this study in the following aspects:
Type of LLMs
In this paper, we only experimented with Llama2-7b-chat, which may limit our understanding of the effectiveness of our approach. A more comprehensive study with different-sized LLMs and prompt formats should be conducted.
Data
Our evaluation was only conducted on two commonly used benchmarks which may limit the diversity in domain, style, and language. There may also be a data contamination problem in the experiment. We believe that more datasets should be used to allow for a more comprehensive evaluation.
Ethics Statement
Our work is built on top of an existing LLM. For this reason, we share the similar potential risks and concerns posed by the underlying LLM. Our method is trained on commonly used training resources of the Machine Translation research community and as such we are not expecting our approach to introduce new areas of risks.
Appendix
Appendix A Related Works
Simultaneous Machine Translation (SimulMT)
is the task to provide real-time translation of a source sentence stream where the goal is to minimize the latency while maximizing the translation quality. A common approach is to train a MT model on prefix-to-prefix dataset to directly predict target tokens based on partial source tokens (Ma et al., 2019). Alternatively, Liu et al. (2020) proposed the incremental decoding framework to leverage the pretrained offline NMT and turn it into a SimulMT model without further training. A core component of SimulMT is a read-write policy to decide at every step whether to wait for another source token (READ) or to generate a target token (WRITE). Previous methods have explored fixed policy, which always waits for $k$ tokens before generation  (Ma et al., 2019; Zhang et al., 2022) and adaptive policy, which trains an agent via reinforcement learning (Gu et al., 2017; Arthur et al., 2021). Re-translation (Arivazhagan et al., 2019) from the beginning of the source sentence at the WRITE step will incur high translation latency. Stable hypothesis detection methods such as Local Agreement, hold-$n$ (Liu et al., 2020) and Share prefix SP-n (Nguyen et al., 2021) are employed to commit stable hypothesis and only regenerate a subsequence of source sentence. The goal is to reduce the latency and minimize the potential for errors resulting from incomplete source sentence (Polák et al., 2022; Wang et al., 2021).
LLM-based NMT
Recent research has delved into the potential usage of LLMs in MT Hendy et al. (2023); Zhu et al. (2023); Robinson et al. (2023), especially in handling discourse phenomena (Wang et al., 2023a; Wu et al., 2024) and linguistic nuances such as idioms (Manakhimova et al., 2023). While LLMs do exhibit some level of translation capability, prior research has identified that they still lags behind the conventional NMT models, especially for low resource languages (Robinson et al., 2023). Additionally, the translation performance varies depending on prompting strategies (Zhang et al., 2023). Efforts have been made to enhance the translation performance of LLMs by incorporating guidance from dictionary (Lu et al., 2023), further fine-tuning (Zeng et al., 2023; Xu et al., 2023) and augmenting with translation memories (Mu et al., 2023).
LLM-based SimulMT
SimulLLM (Agostinelli et al., 2023) explore the ability to adapt an LLM finetuned on NMT task to simultaneous translation with wait-k strategy. Wang et al. (2023b) adopt hybrid READ/WRITE policy with wait-k and incremental decoding. TransLLaMA (Koshkin et al., 2024) teach LLMs to produce WAIT tokens to preserve the causal alignment between source and target tokens. At each inference round, LLMs only produce a single word or WAIT token, which is very costly due to multiple rounds of LLM calls. Guo et al. (2024) introduce LLM into the SimulMT task as a translation agent working with a specialized SimulMT policy agent. An additional memory module stores translation history. The policy agent decides on READ/WRITE actions, while the LLM translates target segments. They face the same KV-cache reuse challenge noted by Wang et al. (2023b), making the computational cost of collaborating big and small models even more significant.
Appendix B Simultaneous Decoding with LLM
0:  LLM : $\texttt{LLM}_{\theta}$, Source chunks: $\mathbf{x}=[]$, Target chunks: $\mathbf{y}=[]$, KV-Cache: $\mathbf{h}=[]$, Chunk index: $c=0$, Variables Definition: Source chunk size: $n$, Beam-size: $B$, Agreement-degree: $\gamma$
1:  while NOT_FINISH do
2:     $\mathbf{x}_{c}\leftarrow\texttt{READ($n$)}$ //READ $n$ tokens
3:     $\mathbf{x}$.append($\mathbf{x}_{c}$)
4:     $\mathbf{x}_{\text{prompt}}\leftarrow\texttt{PROMPT}(\mathbf{x},\mathbf{y})$
[…]
6:     //$B$ candidates with latest tokens in $\mathbf{y}_{c}^{\prime}$
7:     $\mathbf{y}_{c},\mathbf{h}\leftarrow\texttt{PREFIX}(\mathbf{y}_{c}^{\prime},\mathbf{h}^{\prime})$
8:     //Prune with Prefix selection, e.g. RALCP
9:     if $\mathbf{y}_{c}==\emptyset$ then
10:        continue
11:     else
12:        $\mathbf{y}$.append($\mathbf{y}_{c}$)
13:        $\texttt{WRITE}(\mathbf{y}_{c})$
14:        $c\leftarrow c+1$
15:     end if
16:  end while
Algorithm 1 Conversational SimulMT Decoding
Algorithm 1 presents the detail of applying conversational prompt for decoding.

### v4

@@
 Recent studies have shown that LLMs can achieve good performance in SimulMT tasks.
-However, this often comes at the expense of high inference cost and latency.
+However, this often comes at the expense of high inference costs and latency.
 In this paper, we propose a conversational SimulMT framework to enhance the inference efficiency of LLM-based SimulMT through multi-turn-dialogue-based decoding where source and target chunks interleave in translation history, enabling the reuse of Key-Value cache.
 To adapt LLMs to the proposed conversational decoding, we create supervised fine-tuning training data by segmenting parallel sentences using an alignment tool and a novel augmentation technique to enhance generalization.
-Our experiments with Llama2-7b-chat on two SimulMT benchmarks demonstrate that the proposed method empowers the superiority of LLM in translation quality, meanwhile achieves comparable computational latency with specialized SimulMT models.
-[^1] Code, weights, and data will be released with publication.
+Our experiments with Llama2-7b-chat on three SimulMT benchmarks demonstrate that the proposed method empowers the superiority of LLM in translation quality, meanwhile achieving comparable computational latency with specialized SimulMT models.
+[^1] Our code is available at https://github.com/yuriak/LLM-SimulMT
 Conversational SimulMT: Efficient Simultaneous Translation
@@
 1 Introduction
-Simultaneous machine translation (SimulMT) systems provide real-time translation of text input stream Gu et al.
-(2017).
+Simultaneous machine translation (SimulMT) systems provide real-time translation of text input stream gu-etal-2017-learning.
 This task plays an important role in real-world applications, such as facilitating communication in online conferences and generating live subtitles with strict latency requirements.
-In online speech translation, where re-translation is impossible, SimulMT should be robust to errors propagated from the automatic speech recognition system, and produce high-quality translation with low latency.
-(a) Offline Prompt
-(b) Conversational Prompt
-Figure 1: SimulMT prompts in Llama2 format. The system message is blue. The source and corresponding target translations chunks have matching color codes.
-Although large language models (LLMs) have shown the potentials in machine translation (Hendy et al., 2023; Zhu et al., 2023), their applications to SimulMT is non-trivial, as they are not inherently designed for simultaneous decoding.
-Recent works have attempted to adapt LLMs for SimulMT with prefix fine-tuning, incremental decoding (Wang et al., 2023b) and learning to wait for more source tokens before translation (Koshkin et al., 2024).
+Figure 1: Comparison of offline prompt (left) and conversational prompt (right).
+Offline prompt inserts tokens mid-sequence, preventing KV-cache reuse (red X), while conversational prompt appends content sequentially, enabling efficient cache utilization (blue blocks).
+Although large language models (LLMs) have shown the potentials in machine translation (DBLP:journals/corr/abs-2302-09210; DBLP:journals/corr/abs-2304-04675), their applications to SimulMT is non-trivial, as they are not inherently designed for simultaneous decoding.
+Recent works have attempted to adapt LLMs for SimulMT with prefix fine-tuning, incremental decoding (wang2023simultaneous) and learning to wait for more source tokens before translation (koshkin2024transllama).
 These works show LLMs, with careful prompt-engineering, could approach the performance of specialized SimulMT models.
-However, high computational cost, slow inference, and high latency render these approaches impractical for real-world applications (Yuan et al., 2024).
-This is primarily due to the use of offline prompting, where arriving source tokens are inserted at the end of the source sequence, disrupting the continuity of the translation history (Figure 1(a)).
+However, high computational cost, slow inference, and high latency render these approaches impractical for real-world applications (yuan2024llm).
+This is primarily due to the use of offline prompting, where arriving source tokens are inserted at the end of the source sequence, disrupting the continuity of the translation history (Figure˜1 left).
 This prevents reusing cached target history states and requires re-computation of source and target representations.
-To mitigate this issue, we propose conversational prompts that resemble the multi-turn dialogue nature of LLMs.
+To mitigate this issue, we propose conversational prompt that resemble the multi-turn dialogue nature of LLMs.
 Specifically, user inputs are treated as the source tokens to be read, while the LLM’s responses are considered the predicted target tokens to be written.
-In our conversational SimulMT, newly arrived source form the current instruction, while previous source tokens and their translation are treated as conversation history (Figure 1(b)).
-This conversational prompt format enables the reuse of Key-Value cache Pope et al.
-(2023), as all content is appended incrementally without modifying the translation history, hence improving inference efficiency.
-However, conversational SimulMT poses a new challenge for LLMs to accurately identify source sequences, particularly when these sequences are interrupted with target chunks within the conversation history.
-Interleaving incomplete source and target segments in the dialogue history is unnatural (see Figure 1(b)).
-The code-switching phenomenon is exhibited to some extent in LLMs for some languages (Yong et al., 2023); however, it is the continuation rather than the translation of the previous content, making it challenging to leverage existing code-switched datasets for training.
-To adapt the LLM to the conversational decoding format, we take the approach of supervised fine-tuning (SFT) in the conversational SimulMT.
-The challenge is the lack of the conversational SimulMT data for SFT.
-We generate the training data by segmenting parallel sentence pairs into smaller chunks based on a transformation of the word alignments.
-The segmented chunks are further augmented to handle different latency requirements in SimulMT.
-Experiments on two SimulMT benchmarks demonstrate the effectiveness of our proposed conversational SimulMT in balancing the trade-offs between accuracy, speed and flexibility to different latency requirements.
+In our conversational SimulMT, newly arrived source form the current instruction, while previous source tokens and their translations are treated as conversation history (Figure˜1 right).
+This conversational prompt enables the reuse of Key-Value cache kv-cache2023, as all content is appended incrementally without modifying the translation history.
+However, conversational SimulMT poses new challenges for LLMs to comprehend the segmented source content and produce a coherent translation via multi-turn conversation.
+To adapt the LLM to the conversational decoding format, we opt to perform supervised fine-tuning (SFT) on the pretrained LLM.
+But the challenge is the lack of the conversational SimulMT data for SFT.
+Interleaving incomplete source and target segments in the dialogue history is unnatural (see Figure˜1).
+This code-switching style is exhibited in some languages (yong2023prompting); however, it is the continuation rather than the translation of the previous content, making it challenging to leverage existing code-switched datasets for training.
+Therefore, we propose to curate the training data by segmenting parallel sentence pairs into smaller chunks based on a transformation of the word alignments.
+The segmented chunks are further augmented to handle different latency requirements.
+Experiments on three SimulMT benchmarks demonstrate the effectiveness of our proposed conversational SimulMT in balancing the trade-offs between accuracy, speed and flexibility to different latency requirements.
 Compared to offline prompting, our method not only maintains strong performance, but also benefits from reduced latency.
-Notably, our method attains similar decoding speed to the offline LLM-based MT.
+Notably, our method attains similar decoding speed to the LLM-based OfflineMT.
 Our contributions are summarized as follows,
-  • We introduce the conversation prompt to reduce the inference cost of LLM-based SimulMT by leveraging its multi-turn dialogue capability and enabling efficient reuse of Key-Value cached computations.
-  • We present an automated training data curation pipeline which involves a simple heuristic and a novel augmentation strategy to efficiently extract source and target chunks from their word alignments, enabling the LLM to decode with prefix-growing code-switched inputs to support processing of incoming source words while generating target words.
-  • Experiments demonstrate that the proposed conversational SimulMT obtains up to 2$\times$ acceleration compared to the offline-prompting baseline while maintaining comparable translation quality, emphasizing its value in practical applications.
+  • We introduce conversational prompting to reduce the inference cost of LLM-based SimulMT by leveraging its multi-turn dialogue capability and enabling efficient reuse of Key-Value cached computations.
+  • We present an automated training data curation pipeline that can turn any offline translation parallel corpus into the conversational prompt format and generalize with a novel augmentation strategy into any inference setting.
+[…]
 2 Background
 Simultaneous Machine Translation (SimulMT)
-Unlike offline machine translation, where models generate target translation $\mathbf{y}=(y_{1},...,y_{J})$ given a complete source sentence $\mathbf{x}=(x_{1},...,x_{I})$, SimulMT incrementally translates in an online manner with partial source context $\mathbf{x}_{\leq t}=(x_{1},...,x_{t})$ where $I$ and $J$ are source and target length, and $t\leq I$.
-A core component of SimulMT is a read-write policy to decide at every step whether to wait for one or multiple new source tokens (READ) or to generate target tokens given the current partial source sentence (WRITE).
-This policy is crucial in balancing the trade-offs between translation quality and latency.
+Unlike offline machine translation (OfflineMT), where models generate target translation $\mathbf{y}=(y_{1},...,y_{J})$ given a complete source sentence $\mathbf{x}=(x_{1},...,x_{I})$, SimulMT incrementally translates with partial source context $\mathbf{x}_{\leq t}=(x_{1},...,x_{t})$ where $t\leq I$.
+A core component of SimulMT is a read-write policy that decides whether to wait for new source tokens (READ) or generate target tokens (WRITE), balancing translation quality and latency.
 Incremental Decoding
-Studies have explored equipping the capability of simultaneous decoding to an offline neural machine translation model by repetitively performing offline decoding on the incrementally updated source and target histories, without the necessity of READ -WRITE policy learning and model customization (Liu et al., 2020; Nguyen et al., 2021; Polák et al., 2022; Guo et al., 2023).
-Specifically, it involves a chunk-wise READ policy that reads a fixed number of $n$ tokens in each round and a WRITE policy that selects the longest common prefix (LCP) among the beam candidates as a stable hypothesis to be committed Polák et al.
-(2022).
-LCP often results in high latency when no candidates share the same initial token, hindering the WRITE action.
-Relaxed Agreement Longest Common Prefix (RALCP) (Wang et al., 2023b) employs a voting mechanism to accept a common prefix if the agreement among candidates is above an agreement threshold $\gamma$.
-Once a hypothesis is committed, it is no longer modified.
+Studies have explored adapting OfflineMT models for simultaneous decoding by performing offline decoding on incrementally updated histories (low_latency; nguyen21c_interspeech; polak-etal-2022-cuni; HW_TSC).
+This involves a chunk-wise READ policy that reads $n$ tokens per round and a WRITE policy that commits stable partial translations using the longest common prefix (LCP) polak-etal-2022-cuni algorithm.
+LCP often causes high latency when candidates lack common initial tokens.
+Relaxed Agreement LCP (RALCP) (wang2023simultaneous) was proposed to vote for accepting prefixes with candidate agreement above threshold $\gamma$.
 SimulMT with LLMs
-Given that incremental decoding essentially repeats performing offline decoding, leveraging offline-style translation prompts for LLMs is a straightforward approach.
-Meanwhile, offline-style prompts align with LLMs’ innate ability to follow instructions (translating content from one language to another), avoiding the additional fine-tuning Xu et al.
-(2023).
-During each incremental decoding round, a new source chunk is READ in and appended to the source history $\mathbf{x}$.
-The translation hypothesis is WRITTEN out and appended to target history $\mathbf{y}$, forming the latest incremental states.
-Given the new source and target history $\mathbf{x}$ and $\mathbf{y}$, LLMs are prompted to generate the latest translation with the offline prompt as Figure 1(a).
+Since incremental decoding essentially repeats offline decoding, using offline-style translation prompts with LLMs is straightforward and aligns with their instruction-following capabilities xu2023paradigmshift.
+During each round, a source chunk is READ and appended to source history $\mathbf{x}$.
+LLMs generate translations using offline prompts as shown in Figure˜1, which are then WRITTEN to target history $\mathbf{y}$.
 3 Conversational SimulMT
-While incremental decoding with offline prompt enables LLMs to perform simultaneous decoding, it faces high computational latency due to the insertion of newly arrived source tokens in the middle of prompt, disrupting the reuse of cached target history states.
+While incremental decoding with offline prompt enables LLMs to perform simultaneous decoding, it faces high computational latency due to the insertion of newly arrived source tokens in the middle of the prompt, disrupting the reuse of cached target history states.
 In this section, we propose conversational prompts to improve the decoding efficiency and balance quality-latency trade-off.
@@
 the decoding process must consistently add new tokens at the end of the sequence without altering the middle elements.
-When LLMs are performing multi-turn dialogues, the prompt for each turn is composed of a user input and agent response, separated by [INST] and [/INST] tags: “<s>[INST] $X$ [/INST] $Y$</s>", and conversation histories are simply concatenated as the context Touvron et al.
-(2023).
-Drawing parallels to multi-turn dialogues in LLMs, SimulMT can also be viewed similarly, where user inputs are equivalent of READ action, and agent responses are equivalent of WRITE action.
-At round $t$, LLM reads a source context chunk $X_{t}$ and writes its translation $Y_{t}$: “<s>[INST] $X_{t}$ [/INST] $Y_{t}$</s>".
-The conversational prompt of SimulMT can be constructed by placing the source chunk as user input, and the model produces the translation as a response (see Figure 1(b)).
-The conversational prompt enables re-using KV-cache as all contents are appended incrementally.
-Our approach is also able to leverage the hypothesis selection strategy of RALCP (Wang et al., 2023b) to prune the unstable suffixes in each response.
-Algorithm 1 in Appendix B presents the detailed decoding process.
+When LLMs are performing multi-turn dialogues, the prompt for each turn is composed of a user input and assistant response separated by special tokens, and conversation histories are simply concatenated as the context touvron2023llama2.
+Drawing parallels to multi-turn dialogues in LLMs, SimulMT can also be viewed similarly, where user inputs and assistant responses are equivalent to READ and WRITE action.
+At round $t$, LLM reads a source context chunk $X_{t}$ and writes its translation $Y_{t}$: “<U> $X_{t}$ <A> $Y_{t}$".
+The already processed chunks are concatenated as contexts, serving the latest translation round of new incoming chunks.
+As all contents are appended incrementally, the reuse of KV-cache becomes feasible again like in multi-turn dialogue (see Figure˜1).
+Our approach also adapts the hypothesis selection strategy e.g.
+RALCP (wang2023simultaneous) to prune the unstable suffixes in each response.
+Algorithm˜1 in Appendix˜A presents the detailed decoding process.
 Setting N-Shot SacreBLEU COMET
@@
 Reference Flight data showed the plane had to pull out of second
-Table 1: Performance of zero-shot and few-shot conversational SimulMT (Convers.) with Llama2-7b-chat on the WMT15 De->En test set.
-Offline decoding result with the same model is presented (Offline MT) as the baseline.
-A failure case is also shown, in which the LLM tends to complete the sentence in the translation with hallucination (red) leading to a more serious failure in the second chunk.
-We conduct a pilot experiment on the zero-shot and few-shot capability of LLMs in decoding with conversation prompts.
-Specifically, we evaluate the pretrained Llama2-7b-chat (Touvron et al., 2023) with zero- and five-shot settings on the WMT15 De->En test set with chunk size $n$ set as 5.
-As illustrated by Table 1, SimulMT with conversational prompts obtain inferior performance even when instructed with 5-shot prompt.
-The failure case example indicates that the LLM cannot accurately translate the content within the partial source sentence.
-Instead, it tends to complete the sentence in the translation with hallucinated content.
-This is because the LLM is trained on full sentences and, therefore, is not able to handle interleaved partial sentence translation in the multi-turn dialogue format.
-To this end, we propose to perform supervised fine-tuning (SFT) with the LLM on SimulMT data in a conversational format.
-Our approach to tackle the challenge of lacking the conversational SimulMT data for SFT is introduced in the following section.
+Table 1: Performance comparison of Llama2-7b-chat on WMT15 De->En test set in zero-shot and few-shot conversational SimulMT settings.
+OfflineMT results are included as a baseline.
+The example failure case demonstrates how the LLM hallucinates completions (shown in red) when translating partial sentences, leading to compounding errors in subsequent chunks.
+We conducted a pilot experiment to assess LLMs’ zero-shot and few-shot capabilities with conversational prompts.
+Using Llama2-7b-chat (touvron2023llama2) on the WMT15 De->En test set with chunk size $n=5$, we tested both zero- and five-shot settings.
+As shown in Table 1, conversational SimulMT performed poorly even with 5-shot prompting.
+The failure analysis reveals that LLMs, trained primarily on complete sentences, struggle with partial source translation and tend to hallucinate completions when presented with fragments in a multi-turn dialogue format.
+To address this limitation, we propose to SFT LLMs on conversational SimulMT data.
+The following section details our approach to converting a normal bi-text corpus into conversational prompt format.
 Figure 2: The illustration of the data curating process.
 The first graph is obtained from fast_align, it is then modified into a monotonic dependency graph by adding additional edges.
-The Meta Trajectory can be derived by segmenting the monotonic dependency graph with minimal dependency (segment with the dotted line in graph 3).
+The Meta Trajectory can be derived by segmenting the monotonic dependency graph with minimal dependency (segment with the colored solid line in step 3).
 Finally, Policy Generalization is applied to augment the segmented graph with merge (red dotted lines will be removed) and shift (blue dotted lines are shifted) operations.
@@
 As conversational SimulMT data is not naturally available, we propose to synthesize READ / WRITE chunks by segmenting sentence pairs from parallel corpora.
-Inspired by Arthur et al.
-(2021) which generates the oracle policy from word alignments, we further extend the approach by carefully addressing the impact of word reordering and improving the generalizability of the oracle policy.
+Inspired by arthur-etal-2021-learning which generates the oracle policy from word alignments, we further extend the approach by carefully addressing the impact of word reordering and improving the generalizability of the oracle policy.
 Specifically, we first build monotonic dependency graph from the alignment of a sentence pair.
-We then segment the graph and convert these segments into READ / WRITE pairs, followed by augmentation to improve its generalization across various latency demands (Figure 2).
+We then segment the graph and convert these segments into READ / WRITE pairs, followed by augmentation to improve its generalization across various latency demands (Figure˜2).
 The process is explained below.
 Alignment Graph Generation
-Given a sentence pair, we employ fastalign Dyer et al.
-(2013) to obtain word alignment between source and target tokens (Step 1 in Figure 2).
+Given a sentence pair, we employ fastalign fastalign to obtain word alignment between source and target tokens (Step 1 in Figure˜2).
 The obtained alignment is a set $\mathcal{A}$ of pairs $(i,j)$ denoting the source token $x_{i}$ is aligned with its corresponding target token $y_{j}$.
-We define the sufficient source token set to generate a given target token $y_{j}$ as $\textbf{a}_{j}=\{i|(i,j)\in\mathcal{A},\forall i\in[0,I]\}$ of all source tokens needed to produce the target token $y_{j}$.
+We define the sufficient source token set to generate a given target token $y_{j}$ as $\textbf{a}_{j}=\{i|(i,j)\in\mathcal{A},\forall i\in[0,I]\}$.
 A source and target sentences have a monotonic translation relationship if the previous target tokens only aligned with the previous source tokens, i.e.
-$\forall j>k\quad\max(\textbf{a}_{j})\geq\max(\textbf{a}_{k})$ (Koehn et al., 2005; Ling et al., 2011).
+$\forall j>k\quad\min(\textbf{a}_{j})\geq\max(\textbf{a}_{k})$ (koehn-etal-2005-edinburgh; ling-etal-2011-discriminative).
 This condition ensures that the relative order of words is preserved between the source and target sentences.
 In that case, the optimal minimum-latency policy that retains sufficient source information is to produce the monotonic translation that follows the word order of the source sequence, i.e.
-WRITE the target token $y_{j}$ after reading the final source token $x_{\max(\textbf{a}_{j})}$ needed for its translation, and READ the next source tokens.
+WRITE target token $y_{j}$ immediately after reading the final required source token $x_{\max(\textbf{a}_{j})}$, and then READ the next source tokens.
 Monotonic Dependency Graph
-The monotonic dependency between source and target sentences enables the effective implementation of an optimal READ /WRITE policy.
-However, in practice, translations often exhibit non-monotonic relationships, which require re-ordering to produce grammatically correct output in the target language, especially when the source and target languages have significantly different syntactic structures.
-To address this, we propose a heuristic method for constructing a monotonic dependency graph $\overrightarrow{\mathcal{A}}$ from the alignment set $\mathcal{A}$, such that the monotonic condition is met (Step 2 in Figure 2).
-More specifically, for each target token $y_{j}$ which does not satisfy the monotonic condition $\max(\textbf{a}_{j})<\max(\textbf{a}_{j-1})$, we add a new edge from the last sufficient source token $x_{\max(\textbf{a}_{j-1})}$ to $y_{j}$ to eliminate the need of re-ordering.
-For example, in Figure 2, $y_{2}$ violates the monotonic condition as the last source token in its sufficient set $\textbf{a}_{2}=\{1\}$ comes after the last sufficient source token of the previous target $\textbf{a}_{1}=\{1,2\}$.
-Therefore, a new edge from $x_{2}$ to $y_{2}$ is added to the graph.
+Monotonic dependency enables effective implementation of optimal READ /WRITE policies.
+However, translations often require reordering to produce grammatically correct output, especially between languages with different syntactic structures.
+To address this, we propose constructing a monotonic dependency graph $\overrightarrow{\mathcal{A}}$ from alignment set $\mathcal{A}$ (Step 2 in Figure˜2) such that the monotonic condition is met.
+For each target token $y_{j}$ violating the monotonic condition $\min(\textbf{a}_{j})<\max(\textbf{a}_{j-1})$, we add a new edge from the last sufficient source token $x_{\max(\textbf{a}_{j-1})}$ to $y_{j}$, eliminating the need for reordering.
+In Figure˜2, $y_{2}$ violates monotonicity as its earliest required source token $\min(\textbf{a}_{2})=1$ precedes the last required source token for the previous target $\max(\textbf{a}_{1})=2$.
+Thus, we add an edge from $x_{2}$ to $y_{2}$.
 Meta Trajectory
 We then segment the monotonic dependency graph and convert these segments into READ / WRITE pairs, representing the meta trajectory of the oracle policy with minimum latency (Step 3 in Figure 2).
-This segmentation involves examining each target token and identifying its exclusive corresponding source tokens with minimal dependency.
-Each subgraph $\overrightarrow{\mathcal{A}_{j}}$ corresponds with a READ / WRITE pair ($R_{j}$, $W_{j}$) where $W_{j}=\{y_{j}\}$ can be a single target token and $R_{j}=\{x_{i}|i\in\textbf{a}_{j}\setminus\textbf{a}_{j-1}\}$ is a set of new source tokens since the last target token.
-When multiple consecutive target tokens depend on the same source token, we combine their WRITE actions and assign the shared source token to $R_{j}=\{x_{i}\}$, forming $W_{j}=\{y_{j},...,y_{j+n}\}$ with these target tokens.
-This procedure generates a meta READ / WRITE trajectory for the oracle policy, denoted as $RW^{\star}=[(R_{1},W_{1}),...,(R_{C},W_{C})],C\leq I$, representing $C$ chunks.
+We examine each target token to identify its exclusive corresponding source tokens with minimal dependency.
+Each subgraph $\overrightarrow{\mathcal{A}_{j}}$ corresponds to a pair $(R_{j},W_{j})$ where $W_{j}=\{y_{j}\}$ is a target token and $R_{j}=\{x_{i}|i\in\textbf{a}_{j}\setminus\textbf{a}_{j-1}\}$ contains new source tokens required since the previous target.
+When consecutive target tokens depend on the same source token, we combine their WRITE actions, assigning the shared source token to $R_{j}=\{x_{i}\}$ and forming $W_{j}=\{y_{j},...,y_{j+n}\}$.
+This generates a meta trajectory $RW^{\star}=[(R_{1},W_{1}),...,(R_{C},W_{C})],C\leq I$, with $C$ chunks.
 Trajectory Augmentation
@@
 To improve the LLM’s adaptability across various latency demands, we augment the meta-trajectory $RW^{\star}$ with a series of merge and shift operations (Step 4 in Figure 2).
-We first traverse $RW^{\star}$ and randomly merge $\delta$ consecutive READ and WRITE actions, forming new pairs $([R_{c},...,R_{c+\delta}],[W_{c},...,W_{c+\delta}])$, where []delimited-[][$\cdot$][ ⋅ ] is the string concatenation operation.
+We first traverse $RW^{\star}$ and randomly merge $\delta$ consecutive READ and WRITE actions, forming new pairs $([R_{c},...,R_{c+\delta}],[W_{c},...,W_{c+\delta}])$, where $\cdot$] is the string concatenation operation.
 Here, $\delta$ is a variable re-sampled from a uniform distribution $\mathcal{U}(\delta_{\min},\delta_{\max})$ where $\delta_{\min}$ and $\delta_{\max}$ are predefined hyperparameters.
-Additionally, with a probability of $\beta$, we randomly shift a portion of tokens from a WRITE action $W_{c}$ to the next one $W_{c+1}$ in the merged trajectory.
+Additionally, with a probability of $\beta$, we shift a portion of tokens from a WRITE action $W_{c}$ to the next one $W_{c+1}$ in the merged trajectory.
 More specifically, we split $W_{c}$ at a proportion $\rho$ and transfer the latter part to the next pair, resulting in $(R_{c},W_{c}^{<\rho}),(R_{c+1},[W_{c}^{>\rho},W_{c+1}])$ where $\rho$ is sampled from $\mathcal{U}(\rho_{\min},0.9)$ where $\rho_{\min}$ is a hyperparameter.
-The proposed augmentation strategy not only enhances the LLM’s context conditioning ability but also suits the incremental decoding, where the end of the prediction is often truncated by the stable hypothesis selection algorithm.
-The resulting trajectory is a sequence of READ / WRITE chunks where each READ chunk contains source text of differing lengths.
-We then format them in the conversational format to make the trajectory into a text sequence.
-During SFT, we apply cross-entropy loss only on target tokens within WRITE chunks without the shift operation.
+This augmentation enhances the LLM’s context conditioning and suits incremental decoding where prediction endings are often truncated by hypothesis selection algorithms.
+The resulting trajectory consists of READ /WRITE chunks of varying lengths, formatted with conversational prompts for SFT.
+During training, we apply cross-entropy loss only on target tokens within unshifted WRITE chunks.
 4 Experiments
-Trajectory Dimension De-En En-Vi
-Meta-Trajectory #Chunk $10.69\pm 5.51$ $12.98\pm 8.07$
-#SRC word/Chunk $1.74\pm 0.78$ $1.38\pm 0.38$
-#TGT word/Chunk $1.79\pm 0.78$ $1.73\pm 0.49$
-Aug-Trajectory #Chunk $2.74\pm 1.21$ $3.12\pm 1.57$
-#SRC word/Chunk $7.01\pm 3.92$ $5.83\pm 2.76$
-#TGT word/Chunk $7.18\pm 3.90$ $7.35\pm 3.49$
-Table 2: This table presents the statistics of the conversational SimulMT data curated from the training set of both benchmarks, including mean and standard deviation of the number of chunks, number of source/target tokens per chunk in the meta- and augmented-trajectory respectively.
+[…]
+Meta-Trajectory #Chunk $10.69\pm 5.5$ $2.98\pm 8.1 11$$1.94\pm 7.3$
+#SRC word/Chunk $1.74\pm 0.8$ $1.38\pm 0.4$ $1.68\pm 0.5$
+#TGT word/Chunk $1.79\pm 0.8$ $1.73\pm 0.5$ $1.53\pm 0.5$
+Aug-Trajectory #Chunk $2.74\pm 1.2$ $3.12\pm 1.6$ $2.95\pm 1.4$
+#SRC word/Chunk $7.01\pm 3.9$ $5.83\pm 2.8$ $7.02\pm 3.6$
+#TGT word/Chunk $7.18\pm 3.9$ $7.35\pm 3.5$ $6.40\pm 3.2$
+[…]
 4.1 Datasets
 WMT15 De->En
-(4.5M training pairs) We employ newstest2013 (3000 pairs) for validation and utilize newstest2015 (2169 pairs) as our test dataset222www.statmt.org/wmt15/.
-For baseline models, we follow instructions in Zhang et al.
-(2022) to use BPE Sennrich et al.
-(2016) for tokenization with 32,000 merge operations, the vocabulary is shared between German and English.
+(4.5M training pairs) We use newstest2013 (3000 pairs) for validation and newstest2015 (2169 pairs) for testing222www.statmt.org/wmt15/.
 IWSLT15 En->Vi
-(133K training pairs) TED tst2012 (1553 pairs) and tst2013 (1268 pairs) are used as validation and test set333nlp.stanford.edu/projects/nmt/.
-We follow Ma et al.
-(2020); Zhang et al.
-(2022) to preprocess the data for baseline models, resulting in vocabulary sizes of 17K and 7.7K for English and Vietnamese.
-Note that the above tokenization is only applicable for non-LLM baseline models. For LLM-based SimulMT, we use their provided vocabularies and tokenizers.
-Conversational SimulMT Dataset
-We employ fastalign Dyer et al.
-(2013) to obtain word alignment between source and target tokens.
-We randomly merge $\delta_{\min}=2$ to $\delta_{\max}=10$ consecutive segments of the meta trajectories.
-For shift operation, we set the value of $\beta$ and $\rho_{\min}$ to 0.5, i.e.
-with the probability of 0.5, we shift at least 50% portions of tokens in a target segment to the next one.
-Table 2 presents detailed statistical information about the dataset.
-4.2 Evaluation Metrics.
-We use SacreBLEU [^4] BLEU+nrefs:1+case:mixed+eff:no+tok:13a
+(133K training pairs) We employ TED tst2012 (1553 pairs) and tst2013 (1268 pairs) as validation and test sets, respectively333nlp.stanford.edu/projects/nmt/.
+MUST-C En->Zh
+(di-gangi-etal-2019-must) (359k training pairs) This TED talk dataset provides 1349 pairs for validation and the tst-COMMON (2841 pairs) for testing.
+Conversational SimulMT Datasets
+For each dataset, we create conversational prompt versions from their training sets using the approach described in §3.2.
+We employ fastalign fastalign to obtain initial word alignment graphs.
+For trajectory augmentation, we set $\delta_{\min:\max}=(2,10)$ for merging operations.
+For shift operations, both $\beta$ and $\rho_{\min}$ are set to 0.5, meaning we shift at least 50% of tokens in a target segment to the next one with 0.5 probability.
+Table˜2 presents detailed statistics for these datasets.
+4.2 Evaluation Metrics
+We evaluate translation quality and latency using SacreBLEU [^4] BLEU+nrefs:1+case:mixed+eff:no+tok:{13a,zh}
 +smooth:exp+version:2.3.1
- (Post, 2018), COMET [^5] https://huggingface.co/Unbabel/wmt22-cometkiwi-da (Rei et al., 2020) and word-level average lagging (AL) (Ma et al., 2019) to evaluate the quality and latency.
-To measure the computational-aware latency, we use the word wall time (WWT) metric (Wang et al., 2023b), which is the average wall time to predict a word on the same hardware.
+ (matt-2018-sacrebleu), COMET [^5] https://huggingface.co/Unbabel/wmt22-cometkiwi-da (rei2020comet), and word-level average lagging (AL) (ma-etal-2019-stacl).
+To assess computational efficiency, we measure word wall time (WWT) (wang2023simultaneous), which represents the average time required to predict a word on identical hardware.
 4.3 Model Training
-For LLM-based methods, we follow Wang et al.
-(2023b) to utilize Llama2-7b-chat Touvron et al.
-(2023) as the backbone for both language pairs.
-We conduct SFT with QLoRA Hu et al.
-(2022); Dettmers et al.
-(2023) for one epoch to reduce the computation.
-The QLoRA adapters are configured with $r=64$ and $\alpha=16$.
-We set the learning rate to 2e-4, batch size to 48, and employ 4-bit quantization.
-For all experiments, a single A100 GPU is used.
-For offline prompt baselines, we tune the model on the offline bitext corpus.
-The LLMs using conversational prompts are tuned on our curated conversational SimulMT data (§3.2).
+For all LLM-based methods, we use Llama2-7b-chat (touvron2023llama2) as the backbone following wang2023simultaneous.
+We conduct QLoRA-based SFT hu2022lora; dettmers2023qlora for one epoch with $r=64$, $\alpha=16$, learning rate of 2e-4, batch size of 48, and 4-bit quantization on a single A100 GPU.
+Both offline and conversational prompt models are fine-tuned on identical data sources (standard offline style bitext from the aforementioned training sets), but formatted as offline prompts and conversational prompts respectively.
+Figure 3: Translation quality and latency results on three benchmarks.
+Results are presented in three groups with different colors: (i) Encoder-Decoder Transformer baselines (orange), (ii) Offline-Prompt LLMs (blue), and (iii) Conversation-Prompt LLMs (red).
+Offline and Simultaneous decoding are distinguished by the first letter (O/S).
 4.4 Settings
-We compare our proposed conversational SimulMT method against the following baselines
-Offline NMT
-Following (Zhang and Feng, 2022), we train vanilla Transformers Vaswani et al. (2017) of size small (48M) for En->Vi and big (300M) for De->En, and use beam size of 5 during inference.
-SimulMT
-We report the performance of SimulMT with (i) fixed policy Wait-$k$ Ma et al. (2019) and (ii) adaptive policy, including Wait-Info Zhang et al. (2022) and ITST Zhang and Feng (2022).
-  • Wait-$k$ (Ma et al., 2019) propose to achieve simultaneous decoding with an encoder-decoder Transformer by first reading $k$ source tokens and then alternately reading and writing one token per time.
-During decoding, we set $k$ from 1 to 8 for De->En and 4 to 8 for En->Vi.
-  • Wait-Info (Zhang et al., 2022) quantitatively measures the information of each token, ensuring that the information in the target remains below the received source information with the threshold $\mathcal{K}$.
-We set the value of $\mathcal{K}$ from 1 to 8 for both datasets during decoding.
-  • ITST (Zhang and Feng, 2022) calculates the weight of information transferred from each source token to the current target token.
-Based on the total information received, it then determines whether to proceed with the translation of the target token.
-We evaluate with the threshold ranging from 0.1 to 0.7 for both datasets in the experiment.
+We compare our proposed conversational SimulMT against the following baselines:
+Encoder-Decoder Transformers
+We evaluate the performance of a series of specialized Encoder-Decoder Transformer models for both OfflineMT and SimulMT:
+  • Offline NMT: Following (zhang2022itst), we train vanilla Transformer vaswani2017transformer (48M parameters for En->Vi; 300M for De->En and Zh->En) with beam size 5 for inference.
+  • Wait-$k$ (ma-etal-2019-stacl): A fixed policy approach that reads $k$ source tokens before alternating read/write operations.
+We test with $k$ ranging from 1-8 for De->En and Zh->En, 4-8 for En->Vi.
+  • ITST (zhang2022itst) An adaptive policy that measures the information transferred from source to target token and determines when to proceed with translation with a threshold (set as 0.1-0.7 for all datasets).
+  • Wait-Info (zhang-etal-2022-wait) An adaptive policy using token information thresholds ($\mathcal{K}$ from 1-8 for all datasets) to coordinate the timing of translation.
 LLM-based SimulMT
-We compare our conversational prompt against LLM-based SimulMT with offline prompt Wang et al.
-(2023b).
-For both methods, we set $n$=[3,5,7,9,11,13] in the experiment.
-Both approaches utilize RALCP as the stable hypothesis selection method if using beam search decoding (beam=5).
-Additionally, we also report GREEDY selection method (decoding with greedy search without pruning its prediction with prefix selection strategy), which serves as a low-latency baseline.
-LLM-based offline NMT is also reported as the performance upper bound.
-Notations
-For space reason, we abbreviate experimental settings in the following fields: (i) decoding mode: offline (O) or simultaneous (S), (ii) model name: dedicated SimulMT models or LLMs (LM), (iii) prompt type: offline (OP) or conversational prompt (CP) used only for LLMs, (iv) hypothesis selection method (only for LLM-based SimulMT): RALCP or GREEDY.
-For example, our proposed method for simultaneous decoding is written as S:LM:CP:RALCP.
-Figure 3: Translation quality and latency results on two benchmarks.
-For abbreviations, please refer to the Notations description (§4).
-Results are presented in three groups with different colors: (i) specialized MT and SimulMT methods (yellow), (ii) LLM-based offline-prompting SimulMT (blue), and (iii) LLM-based conversational SimulMT (red).
+We compare our conversational prompt approach against the offline prompt method (wang2023simultaneous), using identical READ policies with chunk sizes nn=[3,5,7,9,11,13].
+Both approaches are evaluated with RALCP hypothesis selection (beam=5).
+We also assess greedy decoding (beam=1, no hypothesis selection) with our conversational prompting only (as computational latency baseline), since offline prompting inherently requires hypothesis selection and cannot function with greedy search.
+For reference, we include results from LLM-based OfflineMT as a performance upper bound.
 4.5 Results
-Our preliminary result in Table 1 have shown that out-of-the-box LLMs struggle in zero-shot and few-shot conversational SimulMT.
-In our experiments, we focus on evaluating the impact of fine-tuning LLMs on curated data, with the goal of improving their performance in simultaneous translation using conversational prompts.
-Specifically, we aim to determine if fine-tuning can enable LLMs to handle conversational SimulMT more effectively, and to explore any trade-offs between latency reduction and translation quality.
+Our preliminary study in Table˜1 showed LLMs struggle with zero/few-shot conversational SimulMT.
+Here we examine whether fine-tuning on our curated data enables effective conversational SimulMT, focusing on quality-latency balance.
+Figure 4: Relationship between computational efficiency (Word Wall Time) and translation quality (COMET score) on WMT15 De->En.
+Simultaneous decoding settings are shown as circles, with circle size representing variance across different latency control parameters (e.g.
+$n$).
+Offline settings are represented by diamonds.
+Color coding matches Figure˜3, with our proposed approach highlighted in bold.
 Translation Quality
-Figure 3 shows the BLEU and COMET scores, as well as the average latency on two benchmarks.
-Overall, LLM-base SimulMT (red and blue) outperforms specialized methods (yellow), with the gain up to 3 BLEU scores and 6 COMET scores.
-This suggests that leveraging prior knowledge in LLMs is advantageous for translation.
-Furthermore, given increased latency allowance, LLM-based SimulMT can approach the performance of offline NMT.
-Among LLM-based methods, given the same latency, conversational SimulMT (red) achieves comparable results to its offline-prompting counterparts (blue) in BLEU and even performs better if measured by COMET.
-This shows our approach does not negatively impact the translation capabilities of LLMs.
+As shown in Figure˜3, LLM-based approaches (red and blue) outperform Transformer baselines (yellow) across all language pairs by up to 3 BLEU/10 COMET points.
+With sufficient latency allowance, LLM-based SimulMT even surpasses offline Transformer NMT.
+At equivalent latency levels, our conversational prompting (red) achieves comparable BLEU scores to offline prompting (blue) while often showing better COMET scores.
 Translation Latency
-Compared to offline-promting method (blue), conversational SimulMT (red) can effectively reduce translation latency (1.17 and 1.50 AL in average for De->En and En->Vi) in all latency-range.
-It even reaches the similar latency level of specialized SimulMT (yellow) in En->Vi translation.
-The use of RALCP in conversational SimulMT (S:LM:CP:RALCP) can further boost quality over greedy strategy (S:LM:CP:GREEDY), though at the cost of increased latency due to its more conservative behavior.
-Compared to other baseline approaches, conversational SimulMT obtains a better trade-off between latency and quality, as it fits a wider latency span with more consistent quality.
-More importantly, conversational SimulMT (red) can maintain better quality at the low-latency level (AL<4) compared to specialized models (yellow), making it more superior in practical application where both high quality and low latency are required.
-In contrast, given the same decoding configuration, decoding with offline-prompt (blue) is unable to fit into the low-latency span, making its quality advantageous less competitive compared to specialized approaches (yellow).
+Our conversational SimulMT (red) reduces latency compared to offline prompting (blue), with average reductions of 1.17 and 1.50 AL across all benchmarks.
+For En->Vi and En->Zh, our approach achieves latency comparable to specialized SimulMT models.
+While RALCP (S:LLM-ConvPrompt-RALCP) generally provides better quality than greedy decoding (S:LLM-ConvPrompt-Greedy), the latter offers lower latency.
+Practical Advantages
+Most significantly, our conversational SimulMT (red) maintains superior translation quality at low latency levels (AL<4) compared to specialized models (yellow), making it particularly valuable for practical applications requiring both high quality and low latency.
+In contrast, offline prompting (blue) with identical decoding configurations struggles to operate effectively in the low-latency range, diminishing its quality advantages relative to specialized approaches (yellow).
+These results demonstrate that our conversational prompting approach effectively addresses the efficiency-quality trade-off in simultaneous translation with LLMs.
 5 Analysis
-Figure 4: This figure presents the relationship between computational-aware latency in Word Wall Time and the quality in COMET score.
-For settings of simultaneous decoding, they are represented by circles with the size representing the variance in both metrics under different latency control parameters such as $n$.
-For offline settings, diamond marks are used.
-Colors and notations are consistent with the Figure 3, where our proposed approach is highlighted with bold text.
 5.1 Decoding Speed
-While AL is effective in algorithmically quantifying the delay between the translation and the source input, it does not account for the computational cost associated with the inference process, i.e.
-assuming unit decoding speed across all models.
-In real-world applications, the actual inference wall time is a critical factor that directly impacts user experience.
-A model may exhibit low latency in terms of AL but incur a high inference cost, resulting in substantial time latency.
-To address this, we also report decoding speed using the word wall time (WWT) metric, which measures the real inference time per word (§4.4).
-Detailed result in WMT15 De->En is shown in Figure 4.
-Overall, specialized MT and SimulMT The combination of offline-prompting and RALCP (S:LM:OP:RALCP) is the slowest configuration, significantly lagging behind its offline setting (O:LM:OP), despite of obtaining a good translation quality.
-Conversely, when employing conversational prompting (S:LM:CP:RALCP), it can achieve comparable decoding speed with the offline model (O:LM:CP) while maintaining the translation quality level.
-Moreover, employing greedy hypothesis selection with the conversational prompt (S:LM:CP:GREEDY) achieves a comparable speed to specialized SimulMT (yellow) but with superior quality.
-Figure 5: The performance of the model tested with different levels of trajectory augmentation strategy.
+While Average Lagging (AL) effectively quantifies algorithmic delay between translation and source input, it doesn’t account for computational costs.
+In real-world applications, actual inference time critically impacts user experience: a model with low AL might still deliver poor user experience due to high computational overhead.
+To address this limitation, we evaluate decoding speed using Word Wall Time (WWT), which measures actual inference time per word (§4.4).
+Figure 4 presents detailed WWT results for WMT15 De->En translation.
+Our analysis reveals that offline prompting with RALCP (S:LLM-OffPrompt-RALCP) exhibits the slowest performance, making it impractical despite good translation quality.
+In contrast, our conversational prompting approach with RALCP (S:LLM-ConvPrompt-RALCP) achieves computational efficiency comparable to offline LLM translation (O:LLM-ConvPrompt-Beam=5) while maintaining high translation quality.
+Most notably, our conversational prompting with greedy decoding (S:LLM-ConvPrompt-Greedy) delivers the best efficiency-quality balance—achieving processing speeds comparable to specialized SimulMT models (yellow) while producing significantly better translations.
+These results demonstrate that our approach effectively addresses both algorithmic and computational latency concerns, making it suitable for practical deployment.
+Figure 5: Effect of trajectory augmentation strategies on translation quality (BLEU) and latency (AL) for WMT15 De->En.
+Results compare models trained on meta-trajectories alone versus with merge and shift operations.
 5.2 Effectiveness of Trajectory Augmentation
-We conduct an ablation study to assess the impact of trajectory augmentation methods by supervised fine-tuning LLMs on data derived from (i) meta trajectories only, (ii) incorporating merge operations, and (iii) both merge and shift operations (§3.2).
-The fine-tuning hyper-parameters are kept constant, with only the training data varies.
-Figure 5 reports the performance of conversational SimulMT with RALCP selection strategy under different training data configurations.
-Interestingly, compared to the model trained solely on meta trajectories, trajectory augmentation does not only enhance translation performance but also reduces latency, with the merge operation playing a pivotal role.
-We speculate that without augmentation, the model may have difficulty predicting accurately under varying latency conditions.
-This leads to RACLP accepting less stable hypotheses, hence worsening overall latency.
-5.3 Ability of Context Following
-Figure 6: The model performance for translating the last source chunk w/wo context under different chunk sizes.
-Performing SimulMT with conversational prompts requires the LLM’s proficiency in utilizing context information accurately.
-To further examine the context-following capability of the model tuned with our curated data, we assess its performance in translating the final chunk of the source text, both with and without preceding context.
-Specifically, we extract the inference history for each instance, dividing it into a source-target dialogue history as context and the final source chunk as the latest input.
-Subsequently, the fine-tuned LLM is tasked with generating the last translation chunk from the last source chunk, either with or without the translation history.
-The evaluation is done by computing the BLEU score on the concatenation of the last chunk to its original history.
-A consistent 2-point decrease in performance can be found when decoding without context for all chunk sizes.
-This underscores the LLM’s ability to effectively harness context information for producing more precise translations, as illustrated in Figure 6.
+To evaluate our trajectory augmentation strategy, we conducted an ablation study comparing models trained on: (i) meta trajectories only, (ii) meta trajectories with merge operations, and (iii) meta trajectories with both merge and shift operations (§3.2).
+All models used identical hyperparameters, with training data as the only variable.
+As shown in Figure˜5, trajectory augmentation yields notable improvements in translation quality and latency when using RALCP.
+The merge operation contributes most significantly to these improvements, while models trained solely on meta trajectories perform poorly across all metrics.
+This suggests augmentation techniques enhance the model’s ability to generalize across different latency conditions.
+Without augmentation, the model struggles with varying input chunk sizes, causing RALCP to accept less reliable hypotheses and increasing latency.
+The augmented approach effectively prepares the model for dynamic simultaneous translation scenarios.
+Figure 6: Translation quality (BLEU) on WMT15 De->En when generating the final chunk with vs.
+without preceding context, across different chunk sizes.
+The consistent gap demonstrates effective context utilization.
+5.3 Ability to Leverage Contextual Information
+Effective SimulMT with conversational prompting requires the model’s ability to accurately utilize contextual information.
+To evaluate this capability, we designed an experiment isolating the model’s performance on the final chunk of translation both with and without access to preceding context.
+For each test instance, we extracted the complete inference history and separated it into: (i) the source-target dialogue history serving as context, and (ii) the final source chunk representing the latest input.
+We then tasked our fine-tuned LLM with translating this final chunk under two conditions: with and without access to the preceding conversation history.
+Performance was evaluated by computing BLEU scores on the concatenation of the generated final chunk with its original history.
+As shown in Figure˜6, we observed a consistent 2-point decrease in BLEU scores when context was withheld.
+This performance gap demonstrates our model effectively leverages information from previous conversation turns to produce more accurate translations, confirming the fine-tuned LLM maintains translation coherence.
+5.4 Generalizability Across LLM Families
+(a) Impact of model iteration (Llama-2-7b-chat vs. Llama-3.1-8B-Instruct) on WMT15 De->En.
+(b) Effect of model scale (Llama-3.1-8B-Instruct vs. Llama-3.2-3B-Instruct) on WMT15 De->En.
+(c) Impact of target language proficiency (Llama-3.1-8B-Instruct vs. Qwen2.5-7B-Instruct) on MUST-C En->Zh.
+Figure 7: Performance comparison of different LLM families with our conversational prompt.
+In our main experiments, we used Llama-2-7b-chat following wang2023simultaneous for consistency.
+Now, we examine our approach’s generalizability across different LLMs, using identical training and inference parameters for fair comparison.
+We report only greedy simultaneous decoding and offline beam=5 results to eliminate interference with hypothesis selection.
+Impact of Model Iteration
+We compare Llama-2-7b-chat with the newer Llama-3.1-8B-Instruct (grattafiori2024llama3herdmodels) on WMT15 De->En to assess how model advancements affect performance.
+As shown in Figure˜7(a), the newer model demonstrates consistent improvements in both offline and simultaneous modes.
+This confirms that conversational SimulMT effectively transfers to newer LLMs, with benefits from improved instruction-following capabilities and enhanced language modeling.
+Effect of Model Scale
+We investigate how model size impacts performance by comparing Llama-3.1-8B-Instruct with the smaller Llama-3.2-3B-Instruct (grattafiori2024llama3herdmodels) on WMT15 De->En.
+Figure˜7(b) shows that while the larger model predictably outperforms its smaller counterpart, the 3B model still achieves acceptable translation quality (on par with Llama-2-7b-chat in Figure˜7(a)), suggesting our method is viable on resource-constrained devices.
+Impact of Target Language Proficiency
+We evaluate Llama-3.1-8B-Instruct against Qwen2.5-7B-Instruct (qwen2025qwen25technicalreport) on MUST-C En->Zh to investigate the effect of the model’s target language capabilities.
+As shown in Figure˜7(c), Qwen2.5 consistently outperforms Llama-3.1 for Chinese translation by 1-2 BLEU points across all latency settings, demonstrating that target language proficiency provides additional benefits with our approach.
+6 Related Works
+Simultaneous Machine Translation (SimulMT)
+is the task to provide real-time translation of a source sentence stream where the goal is to minimize the latency while maximizing the translation quality.
+A common approach is to train an MT model on prefix-to-prefix dataset to directly predict target tokens based on partial source tokens (ma-etal-2019-stacl).
+Alternatively, low_latency proposed the incremental decoding framework to leverage the pretrained OfflineNMT model and turn it into a SimulMT model without further training.
+A core component of SimulMT is a read-write policy to decide at every step whether to wait for another source token (READ) or to generate a target token (WRITE).
+Previous methods have explored fixed policy, which always waits for $k$ tokens before generation  (ma-etal-2019-stacl; zhang-etal-2022-wait) and adaptive policy, which trains an agent via reinforcement learning (gu-etal-2017-learning; arthur-etal-2021-learning).
+Re-translation (DBLP:conf/acl/ArivazhaganCMCY19) from the beginning of the source sentence at the WRITE step will incur high translation latency.
+Stable hypothesis detection methods such as Local Agreement, hold-$n$ (low_latency) and Share prefix SP-n (nguyen21c_interspeech) are employed to commit stable hypothesis and only regenerate a subsequence of source sentence.
+The goal is to reduce the latency and minimize the potential for errors resulting from incomplete source sentence (polak-etal-2022-cuni; wang2021diformer).
+LLM-based NMT
+Recent research has delved into the potential usage of LLMs in MT DBLP:journals/corr/abs-2302-09210; DBLP:journals/corr/abs-2304-04675; DBLP:journals/corr/abs-2309-07423, especially in handling discourse phenomena (wang-etal-2023-document-level; wu2024adapting) and linguistic nuances such as idioms (manakhimova-etal-2023-linguistically) and proverbs (wang2025proverbsrunpairsevaluating).
+While LLMs do exhibit some level of translation capability, prior research has identified that they still lags behind the conventional NMT models, especially for low resource languages (DBLP:journals/corr/abs-2309-07423).
+Additionally, the translation performance varies depending on prompting strategies (Zhang2023PromptingLL).
+Efforts have been made to enhance the LLMs’ MT performance by incorporating guidance from dictionary (lu2023chainofdictionary), further fine-tuning (zeng2023tim; xu2023paradigmshift) and augmenting with translation memories (mu-etal-2023-augmenting).
+LLM-based SimulMT
+SimulLLM (agostinelli2023simul) explore the ability to adapt an LLM finetuned on NMT task to simultaneous translation with wait-k strategy.
+wang2023simultaneous adopt hybrid READ/WRITE policy with wait-k and incremental decoding.
+TransLLaMA (koshkin2024transllama) teach LLMs to produce WAIT tokens to preserve the causal alignment between source and target tokens.
+At each inference round, LLMs only produce a single word or WAIT token, which is very costly due to multiple rounds of LLM calls.
+guo2024sillm introduce LLM into the SimulMT task as a translation agent working with a specialized SimulMT policy agent.
+An additional memory module stores translation history.
+The policy agent decides on READ/WRITE actions, while the LLM translates target segments.
+They face the same KV-cache reuse challenge noted by wang2023simultaneous, making the computational cost of collaborating big and small models even more significant.
 7 Conclusion
@@
 We summarize the limitations of this study in the following aspects:
-Type of LLMs
-In this paper, we only experimented with Llama2-7b-chat, which may limit our understanding of the effectiveness of our approach.
-A more comprehensive study with different-sized LLMs and prompt formats should be conducted.
 Data
-Our evaluation was only conducted on two commonly used benchmarks which may limit the diversity in domain, style, and language.
-There may also be a data contamination problem in the experiment.
-We believe that more datasets should be used to allow for a more comprehensive evaluation.
+Our evaluation was conducted on three commonly used benchmarks which may limit the diversity in domains, styles, and languages.
+There may also be potential data contamination concerns since LLMs might have been exposed to parts of our test sets during pre-training.
+A more comprehensive evaluation with diverse datasets across more domains and language pairs would strengthen our findings.
+Alignment-based Data Curation
+Our approach relies on word alignment tools like fast_align to segment parallel sentences, which has inherent limitations.
+These tools may struggle with languages having drastically different word orders or grammatical structures, potentially creating suboptimal segmentation points.
+Furthermore, the alignment quality degrades for distant language pairs or complex sentences with idiomatic expressions and cultural references.
+While our augmentation strategies help mitigate some issues, they are still constrained by the initial alignment quality.
 Ethics Statement
@@
 Appendix
-Appendix A Related Works
-Simultaneous Machine Translation (SimulMT)
-is the task to provide real-time translation of a source sentence stream where the goal is to minimize the latency while maximizing the translation quality.
-A common approach is to train a MT model on prefix-to-prefix dataset to directly predict target tokens based on partial source tokens (Ma et al., 2019).
-Alternatively, Liu et al.
-(2020) proposed the incremental decoding framework to leverage the pretrained offline NMT and turn it into a SimulMT model without further training.
-A core component of SimulMT is a read-write policy to decide at every step whether to wait for another source token (READ) or to generate a target token (WRITE).
-Previous methods have explored fixed policy, which always waits for $k$ tokens before generation  (Ma et al., 2019; Zhang et al., 2022) and adaptive policy, which trains an agent via reinforcement learning (Gu et al., 2017; Arthur et al., 2021).
-Re-translation (Arivazhagan et al., 2019) from the beginning of the source sentence at the WRITE step will incur high translation latency.
-Stable hypothesis detection methods such as Local Agreement, hold-$n$ (Liu et al., 2020) and Share prefix SP-n (Nguyen et al., 2021) are employed to commit stable hypothesis and only regenerate a subsequence of source sentence.
-The goal is to reduce the latency and minimize the potential for errors resulting from incomplete source sentence (Polák et al., 2022; Wang et al., 2021).
-LLM-based NMT
-Recent research has delved into the potential usage of LLMs in MT Hendy et al.
-(2023); Zhu et al.
-(2023); Robinson et al.
-(2023), especially in handling discourse phenomena (Wang et al., 2023a; Wu et al., 2024) and linguistic nuances such as idioms (Manakhimova et al., 2023).
-While LLMs do exhibit some level of translation capability, prior research has identified that they still lags behind the conventional NMT models, especially for low resource languages (Robinson et al., 2023).
-Additionally, the translation performance varies depending on prompting strategies (Zhang et al., 2023).
-Efforts have been made to enhance the translation performance of LLMs by incorporating guidance from dictionary (Lu et al., 2023), further fine-tuning (Zeng et al., 2023; Xu et al., 2023) and augmenting with translation memories (Mu et al., 2023).
-LLM-based SimulMT
-SimulLLM (Agostinelli et al., 2023) explore the ability to adapt an LLM finetuned on NMT task to simultaneous translation with wait-k strategy.
-Wang et al.
-(2023b) adopt hybrid READ/WRITE policy with wait-k and incremental decoding.
-TransLLaMA (Koshkin et al., 2024) teach LLMs to produce WAIT tokens to preserve the causal alignment between source and target tokens.
-At each inference round, LLMs only produce a single word or WAIT token, which is very costly due to multiple rounds of LLM calls.
-Guo et al.
-(2024) introduce LLM into the SimulMT task as a translation agent working with a specialized SimulMT policy agent.
-An additional memory module stores translation history.
-The policy agent decides on READ/WRITE actions, while the LLM translates target segments.
-They face the same KV-cache reuse challenge noted by Wang et al.
-(2023b), making the computational cost of collaborating big and small models even more significant.
-Appendix B Simultaneous Decoding with LLM
+Appendix A Conversational SimulMT Decoding
 0: LLM : $\texttt{LLM}_{\theta}$, Source chunks: $\mathbf{x}=[]$, Target chunks: $\mathbf{y}=[]$, KV-Cache: $\mathbf{h}=[]$, Chunk index: $c=0$, Variables Definition: Source chunk size: $n$, Beam-size: $B$, Agreement-degree: $\gamma$
@@
 4:  $\mathbf{x}_{\text{prompt}}\leftarrow\texttt{PROMPT}(\mathbf{x},\mathbf{y})$
-[…]
+5:  $\mathbf{y}_{c}^{\prime},\mathbf{h}^{\prime}\leftarrow\texttt{LLM}(\mathbf{x}_{\text{prompt}},B,\mathbf{h},\text{latest=True})$
 6:  //$B$ candidates with latest tokens in $\mathbf{y}_{c}^{\prime}$
@@
 Algorithm 1 Conversational SimulMT Decoding
-Algorithm 1 presents the detail of applying conversational prompt for decoding.
+Algorithm˜1 presents the details of applying conversational prompts for decoding.